OpenStudy (anonymous):

A shot-putter throws a 4.00 kg shot at a speed of 10.0 m/s. What is the kinetic energy of the shot? A. 40 J B. 200 J C. 400 J D. 800 J

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large KE = \frac{1}{2}m{v^2} = \frac{1}{2} \times 4 \times {10^2} = ...Joules\]

OpenStudy (anonymous):

200! so 200 J is our solution?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay! thanks! onto the next:)

OpenStudy (michele_laino):

:) ok!